User Tools

Site Tools


The log from experiment vt11c has been copied to:

ANTAB and UVFLG files have been copied to the same place.

Joined this eVLBI experiment later in the schedule.

Calibration scan not done because of wind stow.

40 dB separation between LCP and RCP at 1666 MHz

Agilent 4.2 GHz

Rx Atten 10 dB

LCP only used

DAS profile

Setup: Brett and Ben Lewis

Observers: John D and Slava



lbaops/lbamar2007/vt11cholog.txt · Last modified: 2015/12/18 16:38 by